Find the equation of the line using the given information.
The slope is 3/8 and it passes through the point (x, y) = (1, 4).
_________

Answer by Alan3354(69443)

Similar problem, different numbers, easier.
===================
Find the equation of the line using the point-slope formula. Write the final equation using the slope-intercept form. the x-intercept is 1, and (x, y) (-2,3) is a point on the line
------
x-int is (1,0), a 2nd point.
Slope m = diffy/diffx = 3/-3 = -1
----
y-y1 = m*(x-x1) where (x1,y1) is either point.
y-0 = -1*(x-1)
y = -x + 1